# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

### v1

@@ 8 Data Availability
  • Frozen values. Fixed to the best-fitted values. Calculated in $1-100$ keV energy range. $A_{\odot}$ is the solar iron abundance. $r_{\rm g}=GM/\rm c^{2}$. aKushwaha et al. (2021), bMastroserio et al. (2025), cPeng et al. (2024a).
Table 5: Model-independent polarimetric results obtained using the PCUBE algorithm in the $2-8$ keV energy band from IXPE observations of eleven BH-XRBs. PD and PA represent the polarization degree and angle, respectively. Q///I and U///I are the normalized Stokes parameters. MDP denotes the minimum detectable polarization, and SIGNIF is the detection significance in units of $\sigma$. PDE represents the significance of the energy variation of PD in units of $\sigma$. Gray-shaded entries indicate null-detection of polarization. See text for details.
Source Epoch Obs. ID PD PA Q///I U///I $\rm MDP_{99}$ SIGNIF Spectral Remarks PDE
@@
Cyg X11-1- 1 E1 01002901 $3.70\pm 0.19$ $-20.75\pm 1.50$ $2.77\pm 0.19$ $-2.45\pm 0.19$ $0.59$ $19$ LHS Detection $2.8$
E2 01250101 $3.79\pm 0.32$ $-25.54\pm 2.39$ $2.38\pm 0.32$ $-2.95\pm 0.32$ $0.96$ $12$ LHS Detection $<1$
E3 02008201 $2.50\pm 0.40$ $-19.23\pm 4.63$ $1.95\pm 0.40$ $-1.55\pm 0.40$ $1.22$ $5.7$ HSS Detection $<1$
E4 02008301 $2.46\pm 0.30$ $-22.89\pm 3.45$ $1.71\pm 0.30$ $-1.76\pm 0.30$ $0.90$ $8$ HSS Detection $3$
E5 02008401 $2.01\pm 0.29$ $-25.82\pm 4.11$ $1.25\pm 0.29$ $-1.58\pm 0.29$ $0.88$ $6.6$ HSS Detection $>4$
E6 02008501 $1.44\pm 0.28$ $-25.38\pm 5.66$ $0.91\pm 0.28$ $-1.12\pm 0.28$ $0.86$ $4.5$ HSS Detection $1.2$
E7 02008601 $2.07\pm 0.23$ $-36.37\pm 3.19$ $0.61\pm 0.23$ $-1.98\pm 0.23$ $0.70$ $9$ HSS Detection $1.1$
E8 03002201 $3.75\pm 0.47$ $-24.22\pm 3.61$ $2.49\pm 0.47$ $-2.80\pm 0.47$ $1.43$ $7.6$ LHS Detection $1.8$
E9 03003101 $3.04\pm 0.48$ $-18.35\pm 4.54$ $2.44\pm 0.48$ $-1.82\pm 0.48$ $1.46$ $5.9$ LHS Detection $<1$
E10 03010001 $4.65\pm 0.46$ $-27.77\pm 2.85$ $2.63\pm 0.46$ $-3.83\pm 0.46$ $1.4$ $10.1$ LHS Detection $<1$
E11 03010101 $4.76\pm 0.53$ $-32.64\pm 3.18$ $1.99\pm 0.53$ $-4.32\pm 0.53$ $1.60$ $9$ LHS Detection $1.5$
E12 03002599 $2.79\pm 0.19$ $-21.97\pm 1.97$ $2.01\pm 0.19$ $-1.94\pm 0.19$ $0.58$ $14.5$ HSS Detection $>4$
4U 16304747-47- 47 E1 01250401 $8.34\pm 0.17$ $17.80\pm 0.60$ $6.79\pm 0.17$ $4.86\pm 0.17$ $0.53$ $48$ HSS Detection $>4$
E2 02250601 $6.77\pm 0.21$ $21.36\pm 0.90$ $4.97\pm 0.21$ $4.59\pm 0.21$ $0.65$ $31.7$ SPL Detection $>4$
Cyg X33-3- 3 E1 02001899 $20.60\pm 0.31$ $-89.79\pm 0.43$ $-20.60\pm 0.31$ $-0.15\pm 0.31$ $0.94$ $66.3$ LHS Detection $>4$
E2 02250301 $10.58\pm 0.28$ $-87.39\pm 0.75$ $-10.54\pm 0.28$ $-0.96\pm 0.28$ $0.85$ $38$ HSS Detection $>4$
E3 02009101 $21.41\pm 0.41$ $-88.09\pm 0.55$ $-21.36\pm 0.41$ $-1.43\pm 0.41$ $1.24$ $52.7$ SIMS Detection $>4$
E4 03250301 $11.98\pm 0.43$ $-85.22\pm 1.03$ $-11.82\pm 0.43$ $-1.99\pm 0.43$ $1.31$ $27.8$ HSS Detection $>4$
LMC X11-1- 1 E1 02001901 $1.04\pm 0.40$ $53.97\pm 11.09$ $-0.32\pm 0.40$ $0.99\pm 0.40$ $1.23$ $1.8$ HSS Null-detection --
4U 1957$+115$ E1 02006601 $1.85\pm 0.37$ $-42.09\pm 5.75$ $0.19\pm 0.37$ $-1.84\pm 0.37$ $1.13$ $4.5$ HSS Detection $1.7$
LMC X33-3- 3 E1 02006599 $3.04\pm 0.40$ $-44.24\pm 3.77$ $0.08\pm 0.40$ $-3.04\pm 0.40$ $1.21$ $7.2$ HSS Detection $1.8$
E2 03004899 $2.39\pm 0.37$ $-38.43\pm 4.45$ $0.54\pm 0.37$ $-2.33\pm 0.37$ $1.13$ $6$ HSS Detection $2.3$
E3 03004901 $2.19\pm 0.77$ $-39.42\pm 10.01$ $0.42\pm 0.77$ $-2.15\pm 0.77$ $2.32$ $2.4$ HSS Null-detection --
Swift J1727.816131613-1613- 1613 E1 02250901 $4.72\pm 0.28$ $2.50\pm 1.69$ $4.71\pm 0.28$ $0.41\pm 0.28$ $0.84$ $17$ HIMS Detection $<1$
@@
E4 02251201 $3.81\pm 0.29$ $-0.49\pm 2.22$ $3.81\pm 0.29$ $-0.07\pm 0.29$ $0.89$ $12.8$ SIMS Detection $1.1$
E5 02251301 $3.35\pm 0.32$ $-2.01\pm 2.75$ $3.34\pm 0.32$ $-0.23\pm 0.32$ $0.98$ $10.4$ SIMS Detection $2.5$
E6 03005701 $1.18\pm 0.44$ $-2.74\pm 10.55$ $1.18\pm 0.44$ $-0.11\pm 0.44$ $1.32$ $2$ HSS Null-detection --
E7 03006001 $0.40\pm 0.36$ $6.05\pm 25.83$ $0.39\pm 0.36$ $0.08\pm 0.36$ $1.08$ $<1$ HSS Null-detection --
@@
E2 03005301 $0.47\pm 0.36$ $-25.15\pm 22.07$ $0.30\pm 0.36$ $-0.36\pm 0.36$ $1.09$ $0.8$ HSS Null-detection --
Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 03250201 $0.25\pm 0.26$ $-22.35\pm 29.98$ $0.17\pm 0.26$ $-0.17\pm 0.26$ $0.78$ $0.5$ HSS Null-detection --
IGR J17091362417091362417091-362417091 - 3624 E1111 04250201 $9\pm 1.41$ $83.71\pm 4.48$ $-8.78\pm 1.41$ $1.96\pm 1.41$ $4.27$ $6.1$ LHS Detection $1.2$
MAXI J174429417442941744-2941744 - 294 E1111 04250301 $0.71\pm 0.42$ $-15.37\pm 17.21$ $-0.60\pm 0.42$ $-0.36\pm 0.42$ $1.27$ $1.2$ HSS Null-Detection --
  • Measurements in $3-8$ keV. Null-detection in $2-8$ keV. Computed considering $6$ linear energy bins for all sources except GX $339-4$, 4U 1957$+115$ and IGR J17091362417091362417091-362417091 - 3624, where $4$ linear bins are used.
Table 6: Results from the spectro-polarimetric modeling of IXPE Stokes spectra across different observation epochs of the sources in $2-8$ keV energy band. The parameters, with their standard definitions, are obtained from the constant polarization model (polconst) and the energy-dependent polarization model (polpow). See text for details.
@@
  • aOnly IXPE spectra of DU1 is fitted. Not constrained. bdiskbb is not required. cpowerlaw is not required. dSpectral fitting performed within $3-8$ keV range.

### v2

@@ 8 Data Availability
  • Frozen values. Fixed to the best-fitted values. Calculated in 11001-100 keV energy range. $A_{\odot}$ is the solar iron abundance. $r_{\rm g}=GM/\rm c^{2}$. aKushwaha et al. (2021), bMastroserio et al. (2025), cPeng et al. (2024a).
Table 5: Model-independent polarimetric results obtained using the PCUBE algorithm in the 282-8 keV energy band from IXPE observations of eleven BH-XRBs. PD and PA represent the polarization degree and angle, respectively. Q//I and U//I are the normalized Stokes parameters. MDP denotes the minimum detectable polarization, and SIGNIF is the detection significance in units of $\sigma$. PDE represents the significance of the energy variation of PD in units of $\sigma$. Gray-shaded entries indicate null-detection of polarization. See text for details.
Source Epoch Obs. ID PD PA Q//I U//I $\rm MDP_{99}$ SIGNIF Spectral Remarks PDE
@@
Cyg X1-1 E1 01002901 $3.70\pm 0.19$ $20.75\pm 1.50$ $2.77\pm 0.19$ $2.45\pm 0.19$ $0.59$ 1919 LHS Detection $2.8$
E2 01250101 $3.79\pm 0.32$ $25.54\pm 2.39$ $2.38\pm 0.32$ $2.95\pm 0.32$ $0.96$ $12$ LHS Detection <1<1
E3 02008201 $2.50\pm 0.40$ $19.23\pm 4.63$ $1.95\pm 0.40$ $1.55\pm 0.40$ $1.22$ $5.7$ HSS Detection <1<1
E4 02008301 $46\pm 0.30 22.$$2.89\pm 3.45$ $1.71\pm 0.30$ $1.76\pm 0.30$ $0.90$ 88 HSS Detection 33
E5 02008401 $2.01\pm 0.29$ $25.82\pm 4.11$ $1.25\pm 0.29$ $1.58\pm 0.29$ $0.88$ $6.6$ HSS Detection >4>4
E6 02008501 $1.44\pm 0.28$ $25.38\pm 5.66$ $0.91\pm 0.28$ $1.12\pm 0.28$ $0.86$ $4.5$ HSS Detection $1.2$
E7 02008601 $2.07\pm 0.23$ $36.37\pm 3.19$ $0.61\pm 0.23$ $1.98\pm 0.23$ $0.70$ 99 HSS Detection $1.1$
E8 03002201 $3.75\pm 0.47$ $24.22\pm 3.61$ $2.49\pm 0.47$ $2.80\pm 0.47$ $1.43$ $7.6$ LHS Detection $1.8$
E9 03003101 $3.04\pm 0.48$ $18.35\pm 4.54$ $2.44\pm 0.48$ $1.82\pm 0.48$ $1.46$ $5.9$ LHS Detection <1<1
E10 03010001 $4.65\pm 0.46$ $27.77\pm 2.85$ $2.63\pm 0.46$ $3.83\pm 0.46$ $1.4$ $10.1$ LHS Detection <1<1
E11 03010101 $4.76\pm 0.53$ $32.64\pm 3.18$ $1.99\pm 0.53$ $4.32\pm 0.53$ $1.60$ 99 LHS Detection $1.5$
E12 03002599 $2.79\pm 0.19$ $21.97\pm 1.97$ $2.01\pm 0.19$ $1.94\pm 0.19$ $0.58$ $14.5$ HSS Detection >4>4
4U 163047-47 E1 01250401 $8.34\pm 0.17$ $17.80\pm 0.60$ $6.79\pm 0.17$ $4.86\pm 0.17$ $0.53$ $48$ HSS Detection >4>4
E2 02250601 $6.77\pm 0.21$ $21.36\pm 0.90$ $4.97\pm 0.21$ $4.59\pm 0.21$ $0.65$ $31.7$ SPL Detection >4>4
Cyg X3-3 E1 02001899 $20.60\pm 0.31$ $89.79\pm 0.43$ $20.60\pm 0.31$ $0.15\pm 0.31$ $0.94$ $66.3$ LHS Detection >4>4
E2 02250301 $10.58\pm 0.28$ $87.39\pm 0.75$ $10.54\pm 0.28$ $0.96\pm 0.28$ $0.85$ $38$ HSS Detection >4>4
E3 02009101 $21.41\pm 0.41$ $88.09\pm 0.55$ $21.36\pm 0.41$ $1.43\pm 0.41$ $1.24$ $52.7$ SIMS Detection >4>4
E4 03250301 $11.98\pm 0.43$ $85.22\pm 1.03$ $11.82\pm 0.43$ $1.99\pm 0.43$ $1.31$ $27.8$ HSS Detection >4>4
LMC X1-1 E1 02001901 $1.04\pm 0.40$ $53.97\pm 11.09$ $0.32\pm 0.40$ $0.99\pm 0.40$ $1.23$ $1.8$ HSS Null-detection -
4U 1957+115+115 E1 02006601 $1.85\pm 0.37$ $42.09\pm 5.75$ $0.19\pm 0.37$ $84\pm 0.37 1.1$31.13 $4.5$ HSS Detection $1.7$
LMC X3-3 E1 02006599 $3.04\pm 0.40$ $44.24\pm 3.77$ $0.08\pm 0.40$ $3.04\pm 0.40$ $1.21$ $7.2$ HSS Detection $1.8$
E2 03004899 $2.39\pm 0.37$ $38.43\pm 4.45$ $0.54\pm 0.37$ $2.33\pm 0.37$ $1.13$ 66 HSS Detection $2.3$
E3 03004901 $2.19\pm 0.77$ $39.42\pm 10.01$ $0.42\pm 0.77$ $2.15\pm 0.77$ $2.32$ $2.4$ HSS Null-detection -
E4 03005001 $2.84\pm 0.52$ $47.32\pm 5.24$ $0.23\pm 0.52$ $2.83\pm 0.52$ $1.57$ 55 HSS Detection <1<1
Swift J1727.81613-1613 E1 02250901 $4.72\pm 0.28$ $2.50\pm 1.69$ $4.71\pm 0.28$ $0.41\pm 0.28$ $0.84$ $17$ HIMS Detection <1<1
@@
E4 02251201 $3.81\pm 0.29$ $0.49\pm 2.22$ $3.81\pm 0.29$ $0.07\pm 0.29$ $0.89$ $12.8$ SIMS Detection $1.1$
E5 02251301 $3.35\pm 0.32$ $2.01\pm 2.75$ $3.34\pm 0.32$ $0.23\pm 0.32$ $0.98$ $10.4$ SIMS Detection $2.5$
E6 03005701 $1.18\pm 0.44$ $2.74\pm 10.55$ $1.18\pm 0.44$ $0.11\pm 0.44$ $1.32$ 22 HSS Null-detection -
E7 03006001 $0.40\pm 0.36$ $6.05\pm 25.83$ $39\pm 0.36 0.0$$08\pm 0.36$ $1.08$ <1<1 HSS Null-detection -
@@
E2 03005301 $0.47\pm 0.36$ $25.15\pm 22.07$ $0.30\pm 0.36$ $0.36\pm 0.36$ $1.09$ $0.8$ HSS Null-detection -
Swift J151857.0572147151857.0-572147 E1 03250201 $0.25\pm 0.26$ $22.35\pm 29.98$ $0.17\pm 0.26$ $0.17\pm 0.26$ $0.78$ $0.5$ HSS Null-detection -
IGR J17091362417091-3624 E11 04250201 $9\pm 1.41$ $83.71\pm 4.48$ $8.78\pm 1.41$ $1.96\pm 1.41$ $4.27$ $6.1$ LHS Detection $1.2$
MAXI J17442941744-294 E11 04250301 $0.71\pm 0.42$ $15.37\pm 17.21$ $0.60\pm 0.42$ $0.36\pm 0.42$ $1.27$ $1.2$ HSS Null-Detection -
  • Measurements in 383-8 keV. Null-detection in 282-8 keV. Computed considering $6$ linear energy bins for all sources except GX 3394339-4, 4U 1957+115+115 and IGR J17091362417091-3624, where $4$ linear bins are used.
Table 6: Results from the spectro-polarimetric modeling of IXPE Stokes spectra across different observation epochs of the sources in 282-8 keV energy band. The parameters, with their standard definitions, are obtained from the constant polarization model (polconst) and the energy-dependent polarization model (polpow). See text for details.
@@
  • aOnly IXPE spectra of DU1 is fitted. Not constrained. bdiskbb is not required. cpowerlaw is not required. dSpectral fitting performed within 383-8 keV range.